# Numerical Hopf–Lax formulae for Hamilton–Jacobi equations on unstructured geometries

### v1

@@ 5 Quadratic refinement of the minimum search
One major drawback of the scheme in the form (9) is the relatively low accuracy. We propose therefore a more accurate version, still working without a numerical minimization phase, which will be sketched for simplicity in two space dimensions. In this version, the refinement of the minimum search is performed via a local quadratic approximation.
Let $\mathcal{F}_{j}(y)$ be defined by (14), and denote by $k_{j}$ the index of the node achieving the minimum. Then, we consider the node $x_{k_{j}}$ along with its neighbors (as shown in Fig. 1), and build a least squares quadratic approximation of $\mathcal{F}_{j}$ based on its values on such a stencil.
@@
Finally, this value should replace $\mathcal{F}_{j}(x_{k_{j}})$ in (9). In the case of Dirichlet boundary conditions, if $x_{k_{j}}$ is on the boundary, then no quadratic refinement is obviously performed; if $x_{j}$ is on the boundary, then the updated value $v_{j}^{n+1}$ is obtained as
$v_{j}^{n+1}=\min\left\{b_{j},\min_{y\in\mathbb{R}^{2}}Q(y)\right\}.$
@@
Although we expect the stencil $\{\xi_{l}\}_{1\leq l\leq M}$ to be unisolvent (i.e., that $M\geq 6$) and the matrix $A$ to be positive definite, it might happen that one or both of these conditions fail to hold (for example, a Delaunay triangulation is not necessarily acute). In this case, the simplest countermeasure is to take the node value outright, without applying the quadratic refinement. This leads to a decrease of accuracy only on an expectedly minor portion of the grid.
@@
While this technique proves useful in the time-dependent case (as we will show in the numerical examples), its application to the stationary equation is limited to the value iteration solver, which in some sense stems from a time-dependent formulation. On the contrary, it is not clear how this refinement could be introduced in a policy iteration setting.
@@
The consistency estimate for this version of the scheme will be derived by adapting the arguments used in §2.1, and keeping the simplified 1-D setting. First, note that, given a regular function $u$, the error obtained by a least squares, second-order polynomial approximation of $\mathcal{F}_{j}$ constructed on a fixed local stencils (of radius $O(\Delta x)$ and with a bounded number of nodes), retains the same rate of the interpolation error, i.e., $O(\|\mathcal{F}_{j}^{\prime\prime\prime}\|_{\infty}\Delta x^{3})$. We are not aware of results in this direction (this might be a somewhat classical result), but sketch here the arguments for completeness:
@@ 6 Numerical examples
We present some numerical tests showing both the accuracy and the performance of the proposed schemes for evolutive and stationary HJ equations. Our implementation is written in C language, employing Triangle [11], our favorite easy-to-use library for the generation of Delaunay triangular meshes in two dimensions, and SuiteSparse [10], a sparse matrix library for the solution of linear systems associated to policy improvement and quadratic refinement of the minimum search (although this choice might not be optimal in every respect). Unless otherwise stated, the numerical tests are run on a MacBook Pro equipped with an Intel Core i5 dual-core with 2.9 GHz, 8 Gb RAM and under the OS 11.7.10.
@@ 6.1 Evolutive case
providing classical examples of convex and semiconcave solutions (see [7]). In particular, the first solution exhibits an instantaneous regularization of the initial datum, whereas for the second one, the initial discontinuity of the gradient propagates along expanding circles. We observe that, for this final time $T$, the singularity remains in the interior of the domain.
@@
Finally, we evaluate, under grid refinement, the experimental order of convergence given by $\text{EOC}\sim\log_{2}(\mathcal{E}_{\Delta x}/\mathcal{E}_{\Delta x/2})$, for the errors in both $L^{\infty}$ and $L^{1}$ norms, denoted, respectively, by $\text{EOC}^{\infty}$ and $\text{EOC}^{1}$.
The results for test $1$ and test $2$, associated with the initial conditions $u_{0,1}$ and $u_{0,2}$, are shown in Tables 12, where, for each refinement of the grid, we also report the number of time steps $N$, the sizes $|\mathcal{X}|$ and $|\mathcal{T}|$ of the vertex and triangle lists, and the CPU time of the scheme, expressed in seconds. We set the coefficient $C$ of the displacements (29) to $C=2$.
@@
$\Delta x$ $\Delta t$ $|\mathcal{X}|$ $|\mathcal{T}|$ $\mathcal{E}^{1}_{\Delta x}$ $\text{EOC}^{1}$ $\mathcal{E}^{\infty}_{\Delta x}$ $\text{EOC}^{\infty}$ $N$ CPUtime
0.1 0.1581 3142 6080 0.0523 - 0.0582 - 12 0.41
0.05 0.1118 12360 24329 0.025 1.09 0.031 0.90 17 3.19
0.025 0.0791 49077 97344 0.013 0.9404 0.0153 1.0225 25 27.26
0.0125 0.0559 195420 389229 0.0060 1.0896 0.0068 1.1638 35 210.25
@@
$\Delta x$ $\Delta t$ $|\mathcal{X}|$ $|\mathcal{T}|$ $\mathcal{E}^{1}_{\Delta x}$ $\text{EOC}^{1}$ $\mathcal{E}^{\infty}_{\Delta x}$ $\text{EOC}^{\infty}$ $N$ CPUtime
0.1 0.1581 3142 6080 0.0918 - 0.0917 - 12 0.43
0.05 0.1118 12360 24329 0.0415 1.1451 0.0435 1.0769 17 3.26
0.025 0.0791 49077 97344 0.0198 1.0675 0.0217 1.0024 25 26.31
0.0125 0.0559 195420 389229 0.0094 1.0770 0.01050 1.0478 35 198.99
@@ 6.2 Stationary case
In the second set of examples, we treat the static case, focusing in particular on fast solvers based on PI. As for the evolutive case, we consider the quadratic hamiltonian $H(p)=\frac{1}{2}|p|^{2}$, we fix $\lambda>0$, and we study the following two tests, where we define the function $f$ in (7) as
$f_{3}(x)=\frac{1}{2}(\lambda+1)|x|^{2},\quad f_{4}(x)=\frac{1}{2}(\lambda+1)\min\big{\{}|x-(1,0)|^{2},|x+(1,0)|^{2}\big{\}}.$
@@
$u_{3}(x)=\frac{1}{2}|x|^{2},\quad u_{4}(x)=\frac{1}{2}\min\{|x-(1,0)|^{2},|x+(1,0)|^{2}\},$
for $x\in\Omega\subset\mathbb{R}^{2}$, where $\Omega$ is defined as the circle centered in the origin, with radius 2.
In tables 34, we present the results obtained by applying the value iteration and the policy iteration schemes with rectangular integration to tests 3 and 4, with the exact solutions $u_{3}$ and $u_{4}$. The value iteration starts from a constant initial guess equal to 1, and we set the fictitious time step as $\Delta t=0.5{\Delta x}^{2/3}$, to recover a $2/3$ consistency rate. Tables 56 report the corresponding results for trapezoidal integration, with $\Delta t=0.2\sqrt{\Delta x}$, in order to obtain a unitary consistency rate. We set $\lambda=1$, the displacements length coefficient $C=2$ and the stopping tolerance as $10^{-12}$, which is also used in the internal policy evaluation cycle of the modified PI algorithm.
In the tables, $N$ denotes the number of iterations. The experimental convergence rate follows once more the consistency estimate, while the policy iteration solver results in a significant reduction of the CPU times (about one and a half orders of magnitude for the exact PI), without loss of accuracy in the numerical solution.
0.1 0.1077 2029 3928 0.1524 - 0.1412 -
0.05 0.0678 7919 15555 0.0996 0.6131 0.0910 0.6331
0.025 0.0427 31362 62142 0.0640 0.6381 0.0580 0.6511
0.0125 0.0269 125211 249295 0.0408 0.6505 0.0364 0.6711
$\Delta x$ $N$ CPUtime
0.1 237 4.50
0.05 369 32.78
0.025 574 246.86
0.0125 893 1792.3
(a) Value iteration method
$\Delta x$ $N$ CPUtime
0.1 10 0.18
0.05 15 1.25
0.025 18 7.90
0.0125 24 45.83
(b) Exact PI method
$\Delta x$ $N$ CPUtime
0.1 10 0.39
0.05 15 2.58
0.025 18 16.20
0.0125 24 97.61
(c) Modified PI method
@@
$\Delta x$ $N$ CPUtime
0.1 237 4.53
0.05 369 32.94
0.025 574 240.83
0.0125 893 1818.48
(a) Value iteration method
$\Delta x$ $N$ CPUtime
0.1 12 0.26
0.05 14 1.22
0.025 18 7.69
0.0125 22 40.06
(b) Exact PI method
$\Delta x$ $N$ CPUtime
0.1 12 0.60
0.05 14 2.55
0.025 18 15.788
0.0125 22 87.08
(c) Modified PI method
@@
0.1 0.0632 2029 3928 0.1259 - 0.0940 -
0.05 0.0447 7919 15555 0.0693 0.8624 0.0477 0.9784
0.025 0.0316 31362 62142 0.0361 0.9394 0.0237 1.0116
0.0125 0.0224 125211 249295 0.0191 0.9191 0.0125 0.9182
$\Delta x$ $N$ CPUtime
0.1 394 5.04
0.05 549 33.89
0.025 766 239.34
0.0125 1067 1920.46
(a) Value iteration method
$\Delta x$ $N$ CPUtime
0.1 13 0.17
0.05 16 0.97
0.025 18 5.58
0.0125 23 41.22
(b) Exact PI method
$\Delta x$ $N$ CPUtime
0.1 13 0.36
0.05 16 2.08
0.025 18 12.30
0.0125 23 91.63
(c) Modified PI method
Table 5: Test 3, experimental order of convergence in the case of trapezoidal integration ($\Delta t=0.2\sqrt{\Delta x}$), and comparison of the computational performances associated to the three solvers.
@@
0.1 0.0632 2029 3928 0.1718 - 0.0810 -
0.05 0.0447 7919 15555 0.0975 0.8172 0.0374 1.1152
0.025 0.0316 31362 62142 0.0526 0.8907 0.0193 0.9543
0.0125 0.0224 125211 249295 0.0271 0.9549 0.0104 0.8929
$\Delta x$ $N$ CPUtime
0.1 394 5.09
0.05 549 37.13
0.025 766 262.25
0.0125 1067 1918.26
(a) Value iteration method
$\Delta x$ $N$ CPUtime
0.1 12 0.16
0.05 16 1.23
0.025 20 6.58
0.0125 23 52.25
(b) Exact PI method
$\Delta x$ $N$ CPUtime
0.1 12 0.38
0.05 16 2.66
0.025 20 14.15
0.0125 23 129.48
(c) Modified PI method
@@
6.3 Quadratic refinement
The application of the quadratic refinement previously introduced is studied in the time-dependent case. We repeat test 1 and test 2 adding the minimum search on the quadratic approximation, and we collect the results in Tables 7 and 8, respectively.
We observe that the implementation of this technique results in a relevant reduction of the relative errors, at the cost of an increased CPU time. Figure 5 shows that the increase in accuracy is worth the higher complexity, by comparing the numerical errors in $L^{1}$-norm with respect to the CPU times for the scheme (9) with and without applying quadratic refinement.
Figure 5: Test $1$ (left) and test $2$ (right) relative errors evolution, with respect to the times of implementation. Obtained applying scheme (9) with and without quadratic refinement, respectively for $\Delta x=0.1,0.05,0.025,0.0125$.
Moreover, we observe that the quadratic refinement allows us to recover the quadratic profile of the exact solution near the origin, which is otherwise lost in the previous version of the scheme.
A comparison between the solutions obtained with the two schemes is visually reported in Figure 6, where we present the profile of the two numerical solutions for test 1, where $\Omega$ is defined as a ball of radius 1, obtained sectioning along the plane parallel to the Z-axis passing through the points $(-1,0,0)$ and $(1,0,0)$.
Figure 6: Sections of the numerical solutions of test $1$ at final time $T=2$. The numerical solutions are indicated as $u$QR and $u$ and are obtained applying scheme (9) with and without quadratic refinement, for $\Delta x=0.0125$, $\Delta t=0.5\sqrt{\Delta x}\approx 0.05$.
$\Delta x$ $\Delta t$ $|\mathcal{X}|$ $|\mathcal{T}|$ $\mathcal{E}^{1}_{\Delta x}$ $\text{EOC}^{1}$ $\mathcal{E}^{\infty}_{\Delta x}$ $\text{EOC}^{\infty}$ $N$ CPUtime
0.1 0.1581 3142 6080 0.0112 - 0.0234 - 12 1.34
0.05 0.1118 12360 24329 0.0038 1.5397 0.0088 1.4172 17 8.69
0.025 0.0791 49077 97344 0.0025 0.6030 0.0059 0.5788 25 57.65
0.0125 0.0559 195420 389229 0.0010 1.3899 0.0025 1.2247 35 386.67
Table 7: Test 1 experimental order of convergence for scheme (9) with quadratic refinement application, where $\Delta t=0.5\sqrt{\Delta x}$.
$\Delta x$ $\Delta t$ $|\mathcal{X}|$ $|\mathcal{T}|$ $\mathcal{E}^{1}_{\Delta x}$ $\text{EOC}^{1}$ $\mathcal{E}^{\infty}_{\Delta x}$ $\text{EOC}^{\infty}$ $N$ CPUtime
0.1 0.1581 3142 6080 0.0012 - 0.0118 - 12 1.31
0.05 0.1118 12360 24329 0.0004 1.5727 0.0056 1.0812 17 8.32
0.025 0.0791 49077 97344 0.0002 1.1114 0.0018 1.6059 25 55.68
0.0125 0.0559 195420 389229 0.00008 1.1599 0.0014 0.3370 35 369.05
Table 8: Test 2 experimental order of convergence for scheme (9) with quadratic refinement application, where $\Delta t=0.5\sqrt{\Delta x}$.
6.4 Parallel implementation on GPU and complex geometries

### v2

@@ 5 Quadratic refinement of the minimum search
One major drawback of the scheme in the form (9) is the relatively low accuracy. We propose therefore a more accurate version, still working without a numerical minimization phase, which will be sketched for simplicity in two space dimensions. In this version, the refinement of the minimum search is performed via a local quadratic approximation, which improves accuracy of the scheme at the price of loosing its monotonicity.
Let $\mathcal{F}_{j}(y)$ be defined by (14), and denote by $k_{j}$ the index of the node achieving the minimum. Then, we consider the node $x_{k_{j}}$ along with its neighbours (as shown in Fig. 1), and build a least squares quadratic approximation of $\mathcal{F}_{j}$ based on its values on such a stencil.
@@
Finally, this value should replace $v_{k_{j}}^{n}+\Delta tH^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)$ in (26), whenever $x_{k_{j}}\in\Omega$. If $x_{k_{j}}$ is on the boundary, then we expect that the stencil of quadratic refinement might not be unisolvent; in this case we take the node value instead of performing the refinement.
@@
Although we expect the stencil $\{\xi_{l}\}_{1\leq l\leq M}$ to be unisolvent (i.e., that $M\geq 6$) and the matrix $A$ to be positive definite, it might happen that one or both of these conditions fail to hold. For example, as we have just noted, unisolvency may be lost at the boundary; in our case, this may also occur at internal nodes since the Delaunay triangulation produced by the library Triangle [21] and used in the numerical tests is not necessarily acute. In these cases, the simplest countermeasure is to take the node value outright, without applying the quadratic refinement. Strategies to further improve the accuracy of quadratic refinement with Dirichlet conditions are outside the scope of this paper and will possibly be the object of a future work.
@@
While this technique proves useful in the time-dependent case (as we will show in the numerical examples), its application to the stationary equation is limited to the value iteration solver, which in some sense stems from a time-dependent formulation. On the contrary, it is not clear how this refinement could be introduced in a policy iteration setting. In order to compare the various solvers for the static problem, we will not apply quadratic refinement to this case.
@@
The consistency estimate for this version of the scheme will be derived by adapting the arguments used in §2.1, keeping the simplified 1-D setting and assuming in addition that $H$ has a bounded third derivative. First, note that, given a regular function $u$, the error obtained by a least squares, second-order polynomial approximation of $\mathcal{F}_{j}$ constructed on a fixed local stencils (of radius $O(\Delta x)$ and with a bounded number of nodes), retains the same rate of the interpolation error, i.e., $O(\|\mathcal{F}_{j}^{\prime\prime\prime}\|_{\infty}\Delta x^{3})$. We are not aware of results in this direction (this might be a somewhat classical result), but sketch here the arguments for completeness:
@@ 6 Numerical examples
We present some numerical tests showing both the accuracy and the performance of the proposed schemes for evolutive and stationary HJ equations. Our implementation is written in C language, employing Triangle [21], our favorite easy-to-use library for the generation of Delaunay triangular meshes in two dimensions, and SuiteSparse [20], a sparse matrix library for the solution of linear systems associated to policy evaluation and quadratic refinement of the minimum search (although this choice might not be optimal in every respect). Unless otherwise stated, the numerical tests in serial mode are run on a MacBook Pro equipped with an Intel Core i5 dual-core with 2.9 GHz, 8 Gb RAM under OS 11.7.10, whereas the parallel codes are written in CUDA and running on a GPU server equipped with a Nvidia H100 GPU with 94Gb Ram and 14592 CUDA cores.
@@ 6.1 Evolutive case
providing classical examples of convex and semiconcave solutions (see [9]). In particular, the first solution exhibits an instantaneous regularization of the initial datum, whereas for the second one, the initial discontinuity of the gradient propagates along expanding circles. We observe that, for this final time $T$, the singularity remains in the interior of the domain. Initial conditions and final numerical solutions are shown in Figures 23.
@@
Finally, we evaluate, under grid refinement, the experimental order of convergence given by $\text{EOC}\sim\log_{2}(\mathcal{E}_{\Delta x}/\mathcal{E}_{\Delta x/2})$, for errors in both $L^{\infty}$ and $L^{1}$ norms, denoted, respectively, by $\text{EOC}^{\infty}$ and $\text{EOC}^{1}$.
The results for Test $1$ and Test $2$, associated with the initial conditions $u_{0,1}$ and $u_{0,2}$, are shown in Tables 22, obtained with the coefficient $C$ of the displacements (44) set to $C=2$. In the tables we also report the sizes $|\mathcal{X}|$ and $|\mathcal{T}|$ of the vertex and triangle lists for each refinement of the grid.
@@
0.1 0.1581 3142 6080 0.0523 - 0.0582 -
0.05 0.1118 12360 24329 0.025 1.09 0.031 0.90
0.025 0.0791 49077 97344 0.013 0.9404 0.0153 1.0225
0.0125 0.0559 195420 389229 0.0060 1.0896 0.0068 1.1638
@@
0.1 0.1581 3142 6080 0.0918 - 0.0917 -
0.05 0.1118 12360 24329 0.0415 1.1451 0.0435 1.0769
0.025 0.0791 49077 97344 0.0198 1.0675 0.0217 1.0024
0.0125 0.0559 195420 389229 0.0094 1.0770 0.01050 1.0478
@@
6.2 Quadratic refinement
The application of the quadratic refinement previously introduced is studied in the time-dependent case. We repeat test 1 and test 2 adding the minimum search on the quadratic approximation, and we collect the results in Tables 4 and 4, respectively.
0.1 0.1581 3142 6080 0.0112 - 0.0234 -
0.05 0.1118 12360 24329 0.0038 1.5397 0.0088 1.4172
0.025 0.0791 49077 97344 0.0025 0.6030 0.0059 0.5788
0.0125 0.0559 195420 389229 0.0010 1.3899 0.0025 1.2247
Table 3: Test 1, experimental order of convergence for scheme (9) with quadratic refinement application, where $\Delta t=0.5\sqrt{\Delta x}$.
0.1 0.1581 3142 6080 0.0012 - 0.0118 -
0.05 0.1118 12360 24329 0.0004 1.5727 0.0056 1.0812
0.025 0.0791 49077 97344 0.0002 1.1114 0.0018 1.6059
0.0125 0.0559 195420 389229 0.00008 1.1599 0.0014 0.3370
Table 4: Test 2, experimental order of convergence for scheme (9) with quadratic refinement application, where $\Delta t=0.5\sqrt{\Delta x}$.
We observe that the implementation of this technique results in a relevant reduction of the relative errors, at the cost of an increased CPU time. Figure 4 shows that the increase in accuracy is worth the higher complexity, by comparing the numerical errors in $L^{1}$-norm with respect to the CPU times for the scheme (9) with and without applying quadratic refinement. Note that the improvement is more apparent in Test 2, which has a uniformly semi-concave solution. This agrees with a general drop in the performance of numerical schemes, especially those of higher order, in lack of uniform semi-concavity (see the discussion of this point in [8, 9]).
Figure 4: Test 11 (left) and test 22 (right) relative errors evolution, with respect to CPU times. Obtained applying scheme (9) with and without quadratic refinement, respectively for $\Delta x=0.1,0.05,0.025,0.0125$.
Moreover, we observe that the quadratic refinement allows us to recover the quadratic profile of the exact solution near the origin, which is otherwise lost in the previous version of the scheme. A comparison between the solutions obtained with the two schemes is visually reported in Figure 5, where we present the profile of both numerical solutions for test 1, where $\Omega$ is defined as a ball of radius 1, obtained sectioning along the plane parallel to the Z-axis passing through the points (1,0,0)(-1,0,0) and $(1,0,0)$.
Finally, we note that, despite significantly improving accuracy, this implementation of the quadratic refinement may cause an irregular, and lower than expected, order of convergence, even in the uniform semi-concave case. This might depend on the impossibility to enforce unisolvency (see Remark 3) and is worth a deeper investigation, which is, however, beyond the scope of this paper.
Figure 5: Sections of the numerical solutions of test 11 at final time $T=2$. The numerical solutions are indicated as uuQR and $u$ and are obtained applying scheme (9) with and without quadratic refinement, for $\Delta x=0.0125$, $\Delta t=0.5\sqrt{\Delta x}\approx 0.05$.
@@ 6.3 Stationary case
In the second set of examples, we treat the static case, focusing in particular on the accuracy of the scheme, whereas fast solvers and parallel versions will be treated in the next subsection.
As for the evolutive case, we consider the quadratic Hamiltonian $H(p)=\frac{1}{2}|p|^{2}$, we fix $\lambda>0$, and we study the following two tests for $x\in\Omega\subset\mathbb{R}^{2}$, where $\Omega$ is defined as the circle centered in the origin, with radius 2. In the two cases, we define the function $f$ in (7) as
$f_{3}(x)=\frac{1}{2}(\lambda+1)\min\big\{|x-(1,0)|^{2},|x+(1,0)|^{2}\big\},\qquad f_{4}(x)=3\lambda+\frac{1}{2}-\lambda|x-(1,0)|,$
and enforce the boundary conditions
$b_{3}(x)\equiv 3,\qquad b_{4}(x)=3-\sqrt{5-2x_{1}}.$
@@
$u_{3}(x)=\frac{1}{2}\min\{|x-(1,0)|^{2},|x+(1,0)|^{2}\},\qquad u_{4}(x)=3-|x-(1,0)|,$
which are shown in Fig. 6. Note that the boundary conditions are never active in the former example, while they are attained in the latter.
In tables 88, we present the results obtained by applying the scheme with rectangular integration to tests 3 and 4, with the exact solutions $u_{3}$ and $u_{4}$. The value iteration starts from a constant initial guess equal to 1, and we set the fictitious time step as $\Delta t=0.5{\Delta x}^{2/3}$, to recover a $2/3$ consistency rate. Tables 88 report the corresponding results for trapezoidal integration, with $\Delta t=0.2\sqrt{\Delta x}$, in order to obtain a unitary consistency rate. We set $\lambda=1$, the displacement length coefficient $C=2$ and the stopping tolerance as $10^{-12}$. The experimental convergence rates are in agreement with the consistency estimates.
@@
0.1 0.1077 2318 4274 0.0588 - 0.0769 -
0.05 0.0678 8024 15686 0.0433 0.4777 0.0514 0.5823
0.025 0.0427 31698 62709 0.0268 0.6549 0.0322 0.6734
0.0125 0.0269 125625 249877 0.0172 0.6401 0.0205 0.6544
@@
0.1 0.0632 2029 3928 0.1599 - 0.0754 -
0.05 0.0447 7919 15555 0.0894 0.8384 0.0367 1.0799
0.025 0.0316 31362 62142 0.0471 0.9253 0.0169 1.0748
0.0125 0.0224 125211 249295 0.0233 1.0132 0.0081 1.0623
Table 7: Test 3, experimental order of convergence in the case of trapezoidal integration ($\Delta t=0.2\sqrt{\Delta x}$) and comparison of the computational performances associated to the three solvers.
@@
0.1 0.0632 2318 4274 0.0320 - 0.0496 -
0.05 0.0447 8024 15686 0.0167 0.9409 0.0190 1.3828
0.025 0.0316 31698 62709 0.0102 0.7076 0.0108 0.8111
0.0125 0.0224 125625 249877 0.0046 1.1400 0.0048 1.1777
@@
6.4 Fast solvers, GPU implementation and complex geometries